OpenStudy (anonymous):

Eliminate the parameter. x = 3 cos t, y = 3 sin t Would someone mind helping me with this problem? I don't know how to do this at all

ganeshie8 (ganeshie8):

hint : \(\large \sin^2t + \cos^2t = 1\)

ganeshie8 (ganeshie8):

eliminating parameter simply means cooking up a new equation having only x's and y's

ganeshie8 (ganeshie8):

.... from the given parametric equaitons

OpenStudy (anonymous):

Okay, what do I do with the hint you gave? I'm sorry if I sound like a fool. sin, cos, and tan confuse me terribly

ganeshie8 (ganeshie8):

you only need to remember that identity for doing these problems...

ganeshie8 (ganeshie8):

x = 3 cos t, y = 3 sin t x^2 = ? y^2 = ?

OpenStudy (mathmale):

x = 3 cos t, so x^2=(3cos t)^2=??

OpenStudy (anonymous):

9 cos^2 t??

OpenStudy (mathmale):

Yes, and what about squaring y=3 sin t?

OpenStudy (anonymous):

9 sin^2 t

OpenStudy (mathmale):

Mia: We started with x=3cos t; squaring that produces x^2 = 9 (cos t)^2. Squaring y = 3 sin t produces y^2 = 9 (sin t)^2. Please look at the identity that ganeshie8 shared with us earlier. (sin t)^2 + (cos t)^2 = ? Use this to simplify your expression.

OpenStudy (anonymous):

Why are they being added?

OpenStudy (mathmale):

x^2 + y^2 = 9 (sin t)^2 + 9 (cos t)^2 = 9 ( ?? ) = 9 Good question. The answer revolves around ganeshie8's identity: (sin t)^2 + (cos t)^2 = 1. our job was to eliminate that t, right?

OpenStudy (mathmale):

that x^2 + y^2 (on the far left) equals 9 (on the far right. And x^2 + y^2 = r^2, where r is the radius of the circle in question. We end up with r^2=9. What is r?

OpenStudy (anonymous):

Would that be the 1 from the identity?

OpenStudy (mathmale):

Yes, that's right. r=radius. r^2 = 9 = 3^2. What is r?

OpenStudy (anonymous):

What I meant was, is r=1?

OpenStudy (mathmale):

Oh, I'm sorry. what we get is r-squared on the left and 9, or 3 squared, on the right. Taking the sqrt of both sides, r=3, OK? Think: What does " r " represent?

OpenStudy (anonymous):

You said that r is the radius of the circle

OpenStudy (mathmale):